# Solving Fredholm Integral Equations of the Second Kind via Wasserstein Gradient Flows

### v1

@@ Appendix B Subdifferential of $\mathcal{F}_{\alpha}^{\eta}$
The dominated convergence theorem then guarantees that
$\int\limits_{\mathbb{R}^{d}}\left\langle\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z)\mathrm{d}(t_{n\#}\pi-\pi)(x),t_{n}(z)-z\right\rangle\mathrm{d}\pi(z)\to 0$
as $n\to\infty$.
For the remaining term in (25), using that $\left\|\nabla^{2}\mathrm{k}\right\|\leq\mathtt{M}$ we get that for any $z_{1},z_{2}\in\mathbb{R}^{d}$, $x\in\mathbb{R}^{d}$
$\mathrm{k}(x,z_{2})\geq\mathrm{k}(x,z_{1})+\langle\nabla_{2}\mathrm{k}(x,z_{1}),z_{2}-z_{1}\rangle-\mathtt{M}\left\|z_{1}-z_{2}\right\|^{2}.$
Combing this and [4, Proposition 10.4.2] we get that for any $x\in\mathbb{R}^{d}$ and $\pi\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the strong subdifferential of $\mathcal{L}(\pi,x)$ is given by $\xi_{2}(x,z)=\lambda\nabla_{2}\mathrm{k}(x,z)$, thus
$\lim_{n\to\infty}\{\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\}/\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}\geq 0.$
In addition, since $\mathrm{k}$ is Lipschitz continuous with constant $\mathtt{M}$ we get that for any $\pi_{1},\pi_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x\in\mathbb{R}^{p}$
$|\mathcal{L}(\pi_{1},x)-\mathcal{L}(\pi_{2},x)|\leq\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2}),$
and therefore
$\left|\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right|/\left\|t-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}$ (26)
$\qquad\qquad\qquad\leq\mathtt{M}+\int\limits_{\mathbb{R}^{d}}\left\|\xi_{2}(x,z)\right\|\mathrm{d}\pi(z)$
$\qquad\qquad\qquad\leq\mathtt{M}(1+\lambda).$
@@
$\qquad\qquad\geq-\eta^{-1}\mathtt{M}(1+\lambda)\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}>-\infty.$
In addition, the function
$x\mapsto g^{\prime}(\mathcal{L}(\pi,x))\left(\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right)$
is Lipschitz continuous in $x$, as it is the product of Lipschitz continuous and bounded functions. In particular, $g^{\prime}(\mathcal{L}(\pi,x))$ is upper bounded by $\eta^{-1}$ and the Lipschitz continuity follows since its derivative is bounded (23). Similarly, the function
$x\mapsto\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)$
is upper bounded by (26) and Lipschitz continuous (this follows from the defintion of $\mathcal{L}$ and Assumption 1). Hence, the sequence of functions is equicontinuous and we can apply Fatou’s Lemma for weakly converging measures [24, Theorem 4.1] to obtain
@@
[4, Theorem 10.4.9] shows that the subdifferential of $\mathcal{M}(\pi)=\textup{KL}\left(\pi\middle|\pi_{0}\right)$ is given by
@@
$\partial_{\mathrm{s}}\mathcal{F}_{\alpha}^{\eta}(\pi)$ =$=\left\{x\to-\int\left[\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\pi\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right]\mathrm{d}\pi\left(z\right)\right.$
@@
Appendix C Proof on MKVSDE
C.1 Proof of Proposition 2
To prove existence and uniqueness of the solution of (9) we use standard tools for McKean–Vlasov processes. In particular, we only need to show that the drift of (9), denoted by $b:\mathbb{R}^{d}\times\mathcal{P}_{1}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, given for any $x\in\mathbb{R}^{d}$ and $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ by
$b(x,\pi):=\int\limits_{\mathbb{R}^{d}}b^{\eta}(x,z,\pi)\mathrm{d}\pi(z)-\alpha\nabla U(x),$ (29)
is Lipschitz continuous. The remainder of the proof is classical and is omitted, see for instance [51, Theorem 1.1] or [16, Appendix B.2] for a recent result on a similar scheme.
@@ C.1 Proof of Proposition 2
$\left\|b^{\eta}(x_{1},z,\pi_{1})-b^{\eta}(x_{2},z,\pi_{2})\right\|\leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|),$ (30)
with $\mathtt{C}_{1}$ depending upon $\lambda,\eta$ and $\mathtt{M}$.
Following a similar approach, we can show that $b^{\eta}$ is Lipschitz continuous in $z$: for any $z_{1},z_{2}\in\mathbb{R}^{d}$, $x\in\mathbb{R}^{d}$ and $\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$
$\left\|b^{\eta}(x,z_{1},\nu)-b^{\eta}(x,z_{2},\nu)\right\|$ $\leq\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z_{1},x)}{\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]+\varphi(z_{1})+\eta}-\frac{\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{2},\cdot)\right]+\varphi(z_{2})+\eta}\right\|$
$+\left\|\frac{\lambda\nabla_{1}\mathrm{k}(x,z_{1})+\nabla\varphi(x)}{\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}-\frac{\lambda\nabla_{1}\mathrm{k}(x,z_{2})+\nabla\varphi(x)}{\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right\|$
$\leq\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z_{1},x)-\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]+\varphi(z_{1})+\eta}\right\|$
$+\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]+\varphi(z_{1})+\eta}-\frac{\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{2},\cdot)\right]+\varphi(z_{2})+\eta}\right\|$
$+\lambda(\mathtt{M}/\eta)\left\|z_{1}-z_{2}\right\|$
$\leq 2\lambda(\mathtt{M}/\eta)\left\|z_{1}-z_{2}\right\|$
$+\lambda(\mathtt{M}/\eta^{2})|\lambda\nu\left[\mathrm{k}(z_{2},\cdot)\right]+\varphi(z_{2})-\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]-\varphi(z_{1})|$
$\leq 2\lambda(\mathtt{M}/\eta)\left\|z_{1}-z_{2}\right\|+\lambda(\mathtt{M}^{2}/\eta^{2})(1+\lambda)\left\|z_{1}-z_{2}\right\|$
$\leq\mathtt{C}_{2}\left\|z_{1}-z_{2}\right\|,$
@@
$\qquad\leq\int\left\|b^{\eta}(x_{1},z,\pi_{1})-b^{\eta}(x_{2},z,\pi_{2})\right\|\mathrm{d}\pi_{1}(z)$
$\qquad+|\int b^{\eta}(x_{2},z,\pi_{2})\mathrm{d}(\pi_{1}-\pi_{2})(z)|$
$\qquad\leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|)$
@@
$|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|$ $\leq|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{2},\cdot)\right]|$
@@
$|\pi_{2}\left[b^{\eta}(x_{2},\cdot,\pi_{2})\right]-\pi_{1}\left[b^{\eta}(x_{2},\cdot,\pi_{2})\right]|\leq\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2}).$ (32)
@@
$\qquad\leq\mathtt{C}_{1}((1+\mathtt{M})\left\|x_{1}-x_{2}\right\|+\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2}))+\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2}).$
@@
$\left\|b(x_{1},\pi_{1})-b(x_{2},\pi_{2})\right\|$ $\leq\mathtt{C}_{1}((1+\mathtt{M})\left\|x_{1}-x_{2}\right\|+\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2}))$ (34)
@@
C.2 Proof of Proposition 3
Under Assumption 1, Proposition 1(a) ensures that the functional $\mathcal{F}^{\eta}$ is convex and lower bounded. In addition, Assumption 2 guarantees that $U$ is sufficiently regular to satisfy [30, Assumption 2.2]. To see this, we can use Assumption 2(c) and use the Cauchy-Schwarz inequality to obtain

### v2

@@ Appendix C Subdifferential of $\mathcal{F}_{\alpha}^{\eta}$
The dominated convergence theorem then guarantees that, as $n\to\infty$:
$\int\limits_{\mathbb{R}^{d}}\left\langle\int\limits_{\mathbb{R}^{d}}g^{\prime}(\mathcal{L}(\pi,x))\xi_{2}(x,z)\mathrm{d}(t_{n\#}\pi-\pi)(x),t_{n}(z)-z\right\rangle\mathrm{d}\pi(z)\to 0.$
For the last term in (34), we use the definition of $t_{n}$: as $n\to\infty$ we have
$-\mathtt{C}_{0}\left\|t_{n}(x)-x\right\|^{2}\to 0.$
For the remaining term in (34), using that $\left\|\nabla^{2}\mathrm{k}\right\|\leq\mathtt{M}$ we get that:
$\forall z_{1},z_{2}\in\mathbb{R}^{d},x\in\mathbb{R}^{d}:\qquad\mathrm{k}(x,z_{2})\geq\mathrm{k}(x,z_{1})+\langle\nabla_{2}\mathrm{k}(x,z_{1}),z_{2}-z_{1}\rangle-\mathtt{M}\left\|z_{1}-z_{2}\right\|^{2}.$
Combining this and [4, Proposition 10.4.2] we get that for any $x\in\mathbb{R}^{d}$ and $\pi\in\mathcal{P}_{2}(\mathbb{R}^{d})$, the strong subdifferential of $\mathcal{L}(\pi,x)$ is given by $\xi_{2}(x,z)=\lambda\nabla_{2}\mathrm{k}(x,z)$, thus
$\lim_{n\to\infty}\{\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\}/\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}\geq 0.$ (35)
In addition, since $\mathrm{k}$ is Lipschitz continuous with constant $\mathtt{M}$ we get that for any $\pi_{1},\pi_{2}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ and $x\in\mathbb{R}^{p}$ $|\mathcal{L}(\pi_{1},x)-\mathcal{L}(\pi_{2},x)|\leq\lambda\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2}),$ and therefore
$\left|\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right|/\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}$ (36)
$\qquad\qquad\qquad\leq\lambda\mathtt{M}+\int\limits_{\mathbb{R}^{d}}\left\|\xi_{2}(x,z)\right\|\mathrm{d}\pi(z)\leq 2\lambda\mathtt{M}.$
@@
$\qquad\qquad\geq-\eta^{-1}2\lambda\mathtt{M}\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}>-\infty.$
Let us consider the family of functions
$x\mapsto f_{n}(x):=g^{\prime}(\mathcal{L}(\pi,x))h_{n}(x),$ (37)
where we defined
$h_{n}(x):=\frac{\mathcal{L}(t_{n\#}\pi,x)-\mathcal{L}(\pi,x)-\int\limits_{\mathbb{R}^{d}}\langle\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}.$
By (35), this family satisfies $\liminf_{n\to\infty}f_{n}(x)\geq 0$. Consider the second derivative of (37)
$\nabla f_{n}(x)=g^{\prime}(\mathcal{L}(\pi,x))\nabla h_{n}(x)+h_{n}(x)g^{\prime\prime}(\mathcal{L}(\pi,x))\nabla\mathcal{L}(\pi,x).$
Then, $g^{\prime}(\mathcal{L}(\pi,x))$ is upper bounded by $\eta^{-1}$, $g^{\prime\prime}(\mathcal{L}(\pi,x))$ is upper bounded by $\eta^{-2}$ , $h_{n}$ is upper bounded by (36). Moreover,
$\left\|\nabla\mathcal{L}(\pi,x)\right\|=\left\|\int\limits_{\mathbb{R}^{d}}\xi_{1}(x,z)\mathrm{d}\pi(z)\right\|\leq\mathtt{M}(1+\lambda)$
and, using the Lipschitz continuity of $\nabla_{1}\mathrm{k}$ guaranteed by Assumption 1 and the Cauchy–Schwarz inequality,
$\left\|\nabla h_{n}(x)\right\|$ $=\frac{\left\|\int_{\mathbb{R}^{d}}\nabla_{1}\mathrm{k}(x,z)\mathrm{d}(t_{n\#}\pi-\pi)(z)-\int\limits_{\mathbb{R}^{d}}\langle\nabla_{1}\xi_{2}(x,z),t_{n}(z)-z\rangle\mathrm{d}\pi(z)\right\|}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
$\leq\frac{\int_{\mathbb{R}^{d}}\left\|\nabla_{1}\mathrm{k}(x,t_{n}(z))-\nabla_{1}\mathrm{k}(x,z)\right\|\mathrm{d}\pi(z)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
$+\frac{\int\limits_{\mathbb{R}^{d}}\left\|\nabla_{1}\xi_{2}(x,z)\right\|\left\|t_{n}(z)-z\right\|\mathrm{d}\pi(z)}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
$\leq(1+\lambda)\mathtt{M}\frac{\left(\int_{\mathbb{R}^{d}}\left\|t_{n}(z)-z\right\|^{2}\mathrm{d}\pi(z)\right)^{1/2}}{\left\|t_{n}-\operatorname{Id}\right\|_{\mathrm{L}^{2}(\mathbb{R}^{d},\pi)}}$
$\leq\mathtt{M}(1+\lambda),$
where we used $\xi_{2}(x,z)=\lambda\nabla_{2}\mathrm{k}(x,z)$, Assumption 1 and Jensen’s inequality in the second-to-last inequality.
Hence, the sequence of functions is equicontinuous and we can apply Fatou’s Lemma for weakly converging measures [26, Theorem 4.1] to obtain
@@
[4, Theorem 10.4.9] shows that the subdifferential of $\mathcal{M}(\pi)=\textup{KL}\left(\pi\middle|\pi_{0}\right)$ is
@@
$\partial_{\mathrm{s}}\mathcal{F}_{\alpha}^{\eta}(\pi)$ $=\left\{x\to-\int\left[\frac{\lambda\nabla_{2}\mathrm{k}(z,x)}{\lambda\pi\left[\mathrm{k}(z,\cdot)\right]+\varphi(z)+\eta}+\frac{\lambda\nabla_{1}\mathrm{k}(x,z)+\nabla\varphi(x)}{\lambda\pi\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right]\mathrm{d}\pi\left(z\right)\right$.
@@
Appendix D Proof on MKVSDE
D.1 Proof of Proposition 3
To prove existence and uniqueness of the solution of (9) we use standard tools for McKean–Vlasov processes. In particular, we only need to show that the drift of (9), denoted by $b:\mathbb{R}^{d}\times\mathcal{P}_{1}(\mathbb{R}^{d})\to\mathbb{R}^{d}$, given for any $x\in\mathbb{R}^{d}$ and $\pi\in\mathcal{P}_{1}(\mathbb{R}^{d})$ by (13) is Lipschitz continuous. The remainder of the proof is classical and is omitted, see for instance [53, Theorem 1.1] or [16, Appendix B.2] for a recent result on a similar scheme.
@@ D.1 Proof of Proposition 3
Using the Lipschitz continuity of $\mathrm{k}$ we also have that for any $z\in\mathbb{R}^{d}$
$\hfill|\pi_{2}\left[\mathrm{k}(z,\cdot)\right]-\pi_{1}\left[\mathrm{k}(z,\cdot)\right]|$ $\leq\mathtt{M}\mathbf{W}_{1}(\pi_{1},\pi_{2}).$
@@
$\left\|b^{\eta}(x_{1},z,\pi_{1})-b^{\eta}(x_{2},z,\pi_{2})\right\|$ (40)
$\qquad\qquad\leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|+\mathbf{W}_{1}(\pi_{1},\pi_{2})),$
with $\mathtt{C}_{1}$ depending upon $\lambda,\eta$ and $\mathtt{M}$. Following a similar approach, we can show that $b^{\eta}$ is Lipschitz continuous in $z$: for any $z_{1},z_{2}\in\mathbb{R}^{d}$, $x\in\mathbb{R}^{d}$ and $\nu\in\mathcal{P}_{1}(\mathbb{R}^{d})$
$\left\|b^{\eta}(x,z_{1},\nu)-b^{\eta}(x,z_{2},\nu)\right\|$
$\qquad\qquad\leq\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z_{1},x)}{\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]+\varphi(z_{1})+\eta}-\frac{\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{2},\cdot)\right]+\varphi(z_{2})+\eta}\right\|$
$\qquad\qquad+\left\|\frac{\lambda\nabla_{1}\mathrm{k}(x,z_{1})+\nabla\varphi(x)}{\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}-\frac{\lambda\nabla_{1}\mathrm{k}(x,z_{2})+\nabla\varphi(x)}{\lambda\nu\left[\mathrm{k}(x,\cdot)\right]+\varphi(x)+\eta}\right\|$
$\qquad\qquad\leq\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z_{1},x)-\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]+\varphi(z_{1})+\eta}\right\|$
$\qquad\qquad+\left\|\frac{\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]+\varphi(z_{1})+\eta}-\frac{\lambda\nabla_{2}\mathrm{k}(z_{2},x)}{\lambda\nu\left[\mathrm{k}(z_{2},\cdot)\right]+\varphi(z_{2})+\eta}\right\|$
$\qquad\qquad+\lambda(\mathtt{M}/\eta)\left\|z_{1}-z_{2}\right\|$
$\qquad\qquad\leq 2\lambda(\mathtt{M}/\eta)\left\|z_{1}-z_{2}\right\|$
$\qquad\qquad+\lambda(\mathtt{M}/\eta^{2})|\lambda\nu\left[\mathrm{k}(z_{2},\cdot)\right]+\varphi(z_{2})-\lambda\nu\left[\mathrm{k}(z_{1},\cdot)\right]-\varphi(z_{1})|$
$\qquad\qquad\leq 2\lambda(\mathtt{M}/\eta)\left\|z_{1}-z_{2}\right\|+\lambda(\mathtt{M}^{2}/\eta^{2})(1+\lambda)\left\|z_{1}-z_{2}\right\|$
$\qquad\qquad\leq\mathtt{C}_{2}\left\|z_{1}-z_{2}\right\|,$
@@
$\qquad\leq\int\left\|b^{\eta}(x_{1},z,\pi_{1})-b^{\eta}(x_{2},z,\pi_{2})\right\|\mathrm{d}\pi_{1}(z)+|\int b^{\eta}(x_{2},z,\pi_{2})\mathrm{d}(\pi_{1}-\pi_{2})(z)|$
$\qquad\leq\mathtt{C}_{1}(\left\|x_{1}-x_{2}\right\|+|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|+\mathbf{W}_{1}(\pi_{1},\pi_{2}))$
@@
$\hfill|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{1},\cdot)\right]|$ $\leq|\pi_{2}\left[\mathrm{k}(x_{2},\cdot)\right]-\pi_{1}\left[\mathrm{k}(x_{2},\cdot)\right]|$
@@
$|(\pi_{2}-\pi_{1})\left[b^{\eta}(x_{2},\cdot,\pi_{2})\right]|$ $\leq\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2}).$ (42)
@@
$\qquad\leq\mathtt{C}_{1}(1+\mathtt{M})(\left\|x_{1}-x_{2}\right\|+\mathbf{W}_{1}(\pi_{1},\pi_{2}))+\mathtt{C}_{2}\mathbf{W}_{1}(\pi_{1},\pi_{2}).$
@@
$\left\|b(x_{1},\pi_{1})-b(x_{2},\pi_{2})\right\|$ $\leq\mathtt{C}_{1}(1+\mathtt{M})(\left\|x_{1}-x_{2}\right\|+\mathbf{W}_{1}(\pi_{1},\pi_{2}))$ (44)
@@
D.2 Proof of Proposition 4
Under Assumption 1, Proposition 1(a) ensures that the functional $\mathcal{F}^{\eta}$ is convex and lower bounded. In addition, Assumption 3 guarantees that $U$ is sufficiently regular to satisfy [32, Assumption 2.2]. To see this, we can use Assumption 3(c) and use the Cauchy–Schwarz inequality to obtain